OpenStudy (anonymous):

determine three consecutive odd integers if the square of the largest ineger is 33 less than the sum of the squares of the two smaller integers

myininaya (myininaya):

do you know what consecutive means?

myininaya (myininaya):

it means they come right after each other so if i say n is an odd integer what would be the very next odd integer?

myininaya (myininaya):

ok if 3 is an odd integer and i ask you for the next odd integer after 3, what would you say?

myininaya (myininaya):

hello?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

so it would be n+3,n+5,n+7

OpenStudy (anonymous):

sorry im still new to this site

OpenStudy (radar):

@eliekid are you still here?

OpenStudy (anonymous):

yeah

OpenStudy (radar):

O.K. Welcome to Open Study. Are you an algebra student?

OpenStudy (radar):

We are going to solve for three consecutive odd numbers who have a special relationship to each other. First lets label these odd numbers. Let x = first odd number. Then x +2 = next consecutive odd number Finally x + 4 = the third (and largest) consecutive odd number. Do you see why this is so?

OpenStudy (radar):

I see that you have left so I won't waste my time.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so but my internet is acting funny

OpenStudy (anonymous):

So I got up to (x+4)^2=33-[(x+2)+(x)]

OpenStudy (anonymous):

oh wait nevermind I think I got it now

OpenStudy (anonymous):

actually i got it wrong